INFRARED SPECTRA AND PHOTOCHEMISTRY OF γ-BUTYROLACTONE AND γ-BUTYROLACTONE-d_{6} ISOLATED IN INERT GAS MATRICES

\underline{JONATHAN ROSLUND}, C. A. BAUMANN, Department of Chemistry, University of Scranton, Scranton, PA 18510-4626.

We report the results of ultraviolet photolysis of γ-butyrolactone and γ-butyrolactone-d_{6} trapped in solid nitrogen and argon. The infrared spectra obtained prior to photolysis are in agreement with the gas and liquid phase results obtained by Durig and coworkers^{a} and McDermott^{b}. The carbonyl stretching mode, observed at 1770 cm^{−1} in the liquid, appears as a doublet at 1803 and 1775 cm^{−1} (1797 and 1788 cm^{−1} for the deuterated isotopomer) in the matrix. The photoproducts were characterized via infrared spectroscopy. Irradiation at wavelengths below 250 nm results in the production of CO and CO_{2}. There is also evidence of a secondary photolysis which produces atomic oxygen. Possible mechanisms for the photolysis were evaluated, based on the observed photoproducts and ab initio calculations. These results will be compared with those found for the thermal decomposition of this compound.

aJ. R. Durig, G. L. Coulter, and D. W. Wertz, J. Mol. Spectrosc. 27, 285 (1968).

bD. P. McDermott, J. Phys. Chem. 90, 2569 (1986).